OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 9?

OpenStudy (mikurout):

x^2+y^2 = 9^2

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (x-h)^2+ (y-k)^2=r^2}\) (h,k) is the center r is the radius \(\normalsize\color{blue}{ (x-\color{red}{0} )^2+ (y-\color{red}{0})^2=\color{red}{9}^2}\) \(\normalsize\color{blue}{ x^2+ y^2=81}\)